The mean and the standard deviation of the breaking strength of 100 cables tested by the company is 180 lbs and 25 lbs . Find 95% confidence limit of breaking strength of all cables manufactured by the company



Ans: (180- 24.9 , 180+24.9)


1
Expert's answer
2021-05-07T10:31:38-0400

95%confidence interval ="\bar x±z_{\alpha/2}\frac{s} {\sqrt n}"

The z score corresponding to 95% confidence interval is 1.96.

"=180±1.96×\frac{25}{\sqrt {100}}"

"=180\u00b14.9"

=(175. 1, 184.9)

We are 95% confident that the limit of breaking strength of all cables manufactured by the company is betwee 175.1 and 184.9.